OpenStudy (anonymous):

I REALLY need help on this problem, :/ I've found a whole bunch of different formulas and ended up getting every single option given. >.< May you walk me through this? @ganeshie8 @Directrix

ganeshie8 (ganeshie8):

how many faces it has ?

ganeshie8 (ganeshie8):

top is a circular disk - thats one face

ganeshie8 (ganeshie8):

and the rest of it is a half sphere, right ?

ganeshie8 (ganeshie8):

find the surface area of each face and addup

OpenStudy (anonymous):

1 face. SA of each face!

ganeshie8 (ganeshie8):

Lets calculate the area of top face : Notice that the circle has a radius of `6cm` Area of circle = \(\pi r^2 = \pi (6)^2 = ?\)

OpenStudy (anonymous):

For the Curved Surface Area of Hemisphere : 226. 08 and for Total Surface Area of Hemisphere : 339.12

OpenStudy (anonymous):

Area of circle = 113.1 -All of those are options, that's why I'm confused!

ganeshie8 (ganeshie8):

Excellent !

ganeshie8 (ganeshie8):

You want to cover the entire thing right ?

OpenStudy (anonymous):

Well it says completely, so yes the entire thing..

ganeshie8 (ganeshie8):

So total area of silver foil required = area of circle + area of hemisphere = 113.1 + 226.08 = ?

OpenStudy (anonymous):

= 339. 18 !

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

Great, thanks. I have a few more, do you think you could help me?

ganeshie8 (ganeshie8):

I'll try...